### v1

# Wasserstein Wormhole: Scalable Optimal Transport Distance with Transformers

Abstract
Optimal transport (OT) and the related Wasserstein metric ($W$) are powerful and ubiquitous tools for comparing distributions. However, computing pairwise Wasserstein distances rapidly becomes intractable as cohort size grows. An attractive alternative would be to find an embedding space in which pairwise Euclidean distances map to OT distances, akin to standard multidimensional scaling (MDS). We present Wasserstein Wormhole, a transformer-based autoencoder that embeds empirical distributions into a latent space wherein Euclidean distances approximate OT distances. Extending MDS theory, we show that our objective function implies a bound on the error incurred when embedding non-Euclidean distances. Empirically, distances between Wormhole embeddings closely match Wasserstein distances, enabling linear time computation of OT distances. Along with an encoder that maps distributions to embeddings, Wasserstein Wormhole includes a decoder that maps embeddings back to distributions, allowing for operations in the embedding space to generalize to OT spaces, such as Wasserstein barycenter estimation and OT interpolation. By lending scalability and interpretability to OT approaches, Wasserstein Wormhole unlocks new avenues for data analysis in the fields of computational geometry and single-cell biology.
1 Introduction
Many problems in statistical learning boil down to computing and comparing probability distributions that fit empirical data, and different approaches vary in their attention to distribution shape, spread, and degree of parameterization. Optimal transport methods attempt to find an ideal pairing between individual data points from two distributions [1] by representing distances as the minimum cost of mapping between (weighted) point clouds. While original OT solvers, which rely on linear programming, scale poorly [2], newer approaches can reduce the computational complexity and generalize OT to new domains such as computational biology [3, 4, 5], numerical geometry [6], chemistry [7], and image processing [8].
We are specifically interested in massive cohorts of empirical distributions, each with up to thousands of points. The computational complexity of finding the Wasserstein distance between two point clouds with Sinkhorn’s algorithm [9], the current favored [10] OT solver, is $O(n^{2}\cdot I)$, where $n$ is the size of each point cloud and $I$ is the number of Sinkhorn iterations until convergence. Despite improvements in scalability, current OT algorithms are unable to analyze large cohorts as we must repeat Sinkhorn’s algorithm for all pairs in the cohorts, requiring $O(N^{2}\cdot n^{2}\cdot I)$ time to analyze $N$ distributions.
Here, we present Wasserstein Wormhole, an algorithm that represents each point cloud as a single embedded point, such that the Euclidean distance in the embedding space matches the OT distance between point clouds. In Wormhole space, we compute Euclidean distance in $O(d)$ time for an embedding space with dimension $d$, which acts as an approximate OT distance and enables Wasserstein-based analysis without expensive Sinkhorn iterations.
Point clouds contain a variable number of unordered samples, hence neural architectures that require a fixed-sized or ordered input are not suited to encode them. On the contrary, transformers [11] are uniquely poised to do so, due to their permutation equivariance, indifference to input length, and ability to model relationships between samples in each point cloud via attention.
In addition to an encoder that embeds point clouds, Wormhole trains a decoder that reconstructs point clouds from their embedding. The decoder facilitates interpretability with its ability to explicitly compute Wasserstein barycenters [12] or interpolate between point clouds [13].
Fig. 1: Schematic of Wasserstein Wormhole. Empirical distributions (point clouds) are passed through a transformer to produce per point-cloud vector embeddings such that the Euclidean distance between embeddings match the pairwise Wasserstein distance between point clouds. Since computation of OT distances is laborious, Wormhole is optimized by mini-batches to minimize the discrepancy between the embedding pairwise distances and the pairwise Wasserstein distances of the batch point clouds. The Wormhole decoder (not shown) is a second transformer trained to reproduce the input point clouds from the embedding by minimizing the OT distance between input and output.
The problem solved by Wormhole is analogous to multidimensional scaling [14]. Both methods find an embedding that preserves distances between samples. While several distance metrics are provably capable of being represented in this manner, it has been shown that a host of other ’non-Euclidean’ metrics, including Wasserstein, cannot be exactly equated with Euclidean distance. Counter-intuitively, classical MDS on non-Euclidean distance matrices increases in error as embedding dimensions increase [15].
Inspired by [15], we extend MDS theory to derive upper and lower bounds on the error incurred by our gradient descent-based approach, thereby providing Wormhole with theoretical guarantees on performance for (non-Euclidean) Wasserstein distances. We further describe a projected gradient descent (PGD) [16] algorithm for MDS which is guaranteed to converge to the global optimum for any distance matrix. On small examples where it is feasible to compute bounds and apply PGD, we empirically show that Wormhole approaches the best solution.
2 Prior Work
Learning from point-cloud data is dynamic area of research within machine learning [17], with many Optimal Transport based application. PointNet [18, 19] introduced the idea of permutation-invariant fully-connected neural networks to classify and segment 3D point-clouds, which was later extended into transformer based networks [20]. In the context of OT, several works have noted the expressiveness of Wasserstein space compared to euclidean space [21, 22, 23], and have developed methods to embed point-clouds into Wasserstein space, from which relevant information can be more reliably extracted. These works are in essence the inverse of Wormhole, as their target space is Wasserstein, where in Wormhole, target space is euclidean.
Wasserstein Wormhole belongs to a class of algorithms that exploit structures within distributions or relationships between cohorts of them to accelerate or approximate OT. Assuming the transport plan between two distribution is structured, Low-Rank (LR) Sinkhorn [24] factorizes the transport matrix, forcing it into a low-rank structure which enables faster computation of Sinkhorn iterations. Similarly, Convolutional and Geodesic Sinkhorn [25, 26] construct solvers which leverage the underlying geometry of the point-cloud space to simplify OT optimization. MetaOT [27] recognizes that convergence of Sinkhorn depends greatly on quality of initialization and learns a neural network which learns an initial Sinkhorn potential for a given pair of input distributions. Due to its fully-connected architecture, MetaOT is constrained to distributions which share supports, and can not be applied to arbitrary point-clouds. Although these methods can speed up the computation of an individual OT problem, none alleviate the need for many expensive pairwise computations, and are far slower than Wormhole on large cohorts, as shown in Figure 2.
Like Wormhole, methods such as Deep Wasserstein Embeddings (DWE) [28] and DiffusionEMD [29] produce embeddings that preserve Wasserstein distances. DWE learns a neural network via a Siamese architecture trained to match distances, and DiffusionEMD extends the wavelet-based approximation of EMD [30] onto graph domains. Unlike our Wormhole approach, which applies to point clouds in any dimension, DWE is constrained to small 2D images or 3D grids, while scalability of DiffusionEMD is limited to distributions that share supports, such as grayscale images or voxelized spaces. We benchmark Wormhole against DiffusionEMD & DWE, and show that our method is superior at matching OT distances and deriving insights from embeddings, and Wormhole is the only embedding method which scales to high-dimensional datasets (Table 1).
Fig. 2: Benchmarking run-time of Wormhole against other OT acceleration algorithms. Sampling cohorts of different sizes from the MNIST dataset, we measured the time required for current acceleration algorithms (with GPU implementation in JAX-OTT) and Wormhole to compute or approximate the pairwise Wasserstein matrix. Other methods are more appropriate at tiny cohorts, as they do not require training a parametric model. However, even in cohorts with relatively few samples, Wormhole is superior and no other method can scale to complete datasets, requiring weeks of compute time, even on a fully-utilized 80GB GPU.
Wormhole follows a long line of neural network approaches for Optimal Transport estimation [31]. Sinkhorn iterations, which are quadratic with size of point-clouds, cannot be applied when When point-clouds are massive consisting of thousands of points. Neural OT [32] parameterizes the transport plan between two empirical distributions as a input-convex neural network [33] which allows for fast, mini-batch training. Such models have been incredibly successful at extending OT to scRNA-seq data [34, 5, 35], where samples commonly contain over $10^{5}$ cells. Wormhole works best in a different setting, where individual OT problems can be solved, but their solvers are not efficient enough to scale to the many pairwise comparisons within a large cohort. Future extensions of our work can make use of neural solvers to unlock Wasserstein analysis to large cohorts of massive point-clouds.
3 Preliminaries
3.1 Optimal Transport
We follow the formulation of optimal transport presented in [2]. Given two discrete distributions $P_{x}=\sum_{i=1}^{n}\mu_{i}\delta_{x_{i}}$ and $P_{y}=\sum_{i=1}^{m}\nu_{i}\delta_{y_{i}}$ ($\delta$ is the Dirac function), their entropically regularized OT distance is:
$W_{\varepsilon}(P_{x},P_{y})=\min_{P\in U(\nu,\mu)}\sum_{i,j}c(x_{i},y_{j})P_{i,j}\ -\varepsilon H(P)$ (1)
Where $U(\nu,\mu):=\{P\in R_{+}^{n,m}:\sum_{j}P_{i,j}=\nu_{i},\sum_{i}P_{i,j}=\mu_{j}\}$ is the set of all valid pairing matrices, $c(x_{i},y_{j})$ is the ground cost function between points in $P_{x}$ and $P_{y}$, usually chosen to be the Euclidean distance, and $H(P)=-\sum_{i,j}P_{i,j}(\log P_{i,j}-1)$ is the entropic term. While the unregularized problem incurs a heavy computational burden for distributions with large supports or in high dimensions, we can compute $W_{\varepsilon}$ via a series of GPU-accelerable vector-matrix products known as Sinkhorn iterations [9].
Due to the regularization term, the $W_{\varepsilon}$ distance between a distribution and itself is not 0, which can introduce biases when used to optimize generative models. To allow for effective optimization while retaining the computational advantages of entropic regularization, a common alternative is the Sinkhorn Divergence [36], which subtracts the self-transport costs from $W_{\varepsilon}$:
$S_{\varepsilon}(P_{x},P_{y})=W_{\varepsilon}(P_{x},P_{y})-\frac{W_{\varepsilon}(P_{x},P_{x})+W_{\varepsilon}(P_{y},P_{y})}{2}$ (2)
3.2 Multi-dimensional Scaling
Given the pairwise Wasserstein distance matrix between a cohort of distributions $D_{i,j}=W_{\varepsilon}(X_{i},X_{j})$, we seek an embedding per distribution, such that pairwise Euclidean distance between embeddings match the OT distance $\lVert\alpha_{i}-\alpha_{j}\rVert_{2}^{2}\approx W_{\varepsilon}(X_{i},X_{j}))$. In essence, this describes the MDS problem with the critical difference that for large cohorts, it is computationally infeasible to compute the OT distance between all pairs of distributions. Still, MDS can provide key theoretical insights for our setting.
Definition 3.1 (Euclidean distance matrix (EDM)).
An $N\times N$ distance matrix $D$ is Euclidean if there exists a set of points $X=\{x_{1},x_{2},\dots,x_{N}\}$ in some dimension $x_{i}\in R^{d}$ such that:
$\forall{i,j}:\;D_{i,j}=\|x_{i}-x_{j}\|_{2}^{2}$ (3)
From [37], a matrix $D$ is Euclidean if and only if all of its values are non-negative, it is hollow (diagonals are 0) and the criterion matrix $F$ is positive-semidefinite (PSD):
$F=-JDJ$ (4)
Where $J=I_{N}-\frac{\mathbf{1}_{N}\mathbf{1}_{N}^{\top}}{N}$ is the centering matrix and $1_{N}$ is a vector of length $N$ filled with ones. From brevity, we will call matrices which fit these criterion Euclidean.
Wasserstein distance is not Euclidean (see section 8.3 in [2] and [38]), and there are many examples of cohorts of distributions (Figures S1,S2) wherein pairwise OT distance matrices do not fit the criteria in equation 4. Unlike the Euclidean scenario where classical MDS (cMDS) improves as embedding size grows, on non-Euclidean distance matrices the error behaves counter-intuitively and worsens as dimension increases [15]. This motivates us to seek a description of the theory underpinning the embeddings of non-Euclidean distances.
Dataset Parameters OT Correlation Label Accuracy
Name Cohort Size Cloud Size Dimension Wormhole DiffusionEMD DWE Fréchet Wormhole DiffusionEMD DWE Fréchet
MNIST 70,000 105 2 0.986 0.845 0.923 0.862 0.98 0.84 0.97 0.49
Fashion-MNIST 70,000 356 2 0.999 0.971 0.985 0.992 0.87 0.77 0.87 0.73
ModelNet40 12,311 2048 3 0.996 0.852 0.979 0.952 0.86 0.69 0.78 0.61
ShapeNet 15,011 2048 3 0.995 0.817 0.977 0.973 0.98 0.94 0.97 0.92
Rotated ShapeNet (GW) 15,011 512 3 0.987 NA NA 0.687 0.82 NA NA 0.16
MERFISH Cell Niches 256,058 11 254 0.977 OOM OOM OOM 0.98 OOM OOM OOM
SeqFISH Cell Niches 57,407 29 351 0.984 OOM OOM OOM 0.97 OOM OOM OOM
scRNA-seq Atlas 2,185 69 2500 0.986 OOM OOM OOM 0.96 OOM OOM OOM
Table 1: Summary of each dataset used in this study and results from every benchmarked algorithm. Cloud size denotes the median number of points in each point cloud for every dataset. Pearson correlations are computed between true and estimated Wasserstein distances according to pairwise matrices of $10$ instances of $128$ randomly chosen point clouds from each dataset (OT Correlation). The mean squared error (MSE) follows the same trend with Wormhole outperforming competing methods (Table 3). Object classification accuracy is based on the embeddings of each algorithm. After computing encodings, a classifier was trained on the training set embeddings and we denote the classifier’s accuracy on the test set (Label Accuracy). Rotated ShapeNet is the ShapeNet dataset where each sample is randomly rotated in 3D, subsampled by a factor of $4$, and embedded via Gromov-Wasserstein (GW). We divide the results into Wasserstein on 2D and 3D point-clouds, Gromov-Wasserstein on rotated data with arbitrary orientation and high-dimensional datasets. Full implementation details are described in the supplement (Section 11). NA, Not Applicable. OOM, Out-Of-Memory.
4 Wasserstein Wormhole
Algorithm 1 Wasserstein Wormhole At each training step, a mini-batch of point clouds is randomly sampled from the cohort. Wormhole is optimized so that pairwise distance between mini-batch embeddings match their OT distance and output decodings reconstruct input point clouds.
  Input: point clouds $\Omega=\{X_{i}\}^{N}_{i=1}$, initialized encoder $T$ and decoder $G$, batch size $B$, learning rate $\varepsilon_{t}$
  repeat
     Sample $B$ point clouds $\{x_{i}\}_{i=1}^{B}$ from $\Omega$
     Calculate encodings $\alpha_{i}=T(x_{i})$
     Compute pairwise OT in batch $D_{i,j}=S_{\varepsilon}(x_{i},x_{j})$
     Evaluate stress $\mathcal{L}_{enc}=\sum_{i,j}(\lVert\alpha_{i}-\alpha_{j}\rVert_{2}^{2}-D_{i,j})^{2}$
     Produce decodings $\hat{x}_{i}=G(\alpha_{i})$
     Evaluate decoding error $\mathcal{L}_{dec}=\sum_{i}S_{\varepsilon}(x_{i},\hat{x}_{i})$
     Update encoder $T\leftarrow T-\varepsilon_{t}\nabla(\mathcal{L}_{enc}+\mathcal{L}_{dec})$
     Update decoder $G\leftarrow G-\varepsilon_{t}\nabla(\mathcal{L}_{dec})$
  until Convergence
The complexity of modern OT solvers is quadratic with respect to the number of points in a point cloud; thus, it is only straightfoward to compute pairwise Wasserstein distance for small cohorts of empirical distributions. Indeed, naive approaches will not even scale to standard datasets, such as the 12,311-point-cloud ModelNet40 benchmark [39].
This problem calls for a stochastic approach. Instead of directly assigning each point cloud with an embedding, we propose learning a parametric encoding function. Formally, the dataset in our setting is a cohort of point clouds $\Omega=\{X_{1},X_{2},X_{3},\dots X_{N}\}$, where each point cloud $X_{i}$ contains $n_{i}$ points in $d$ dimensions. Our goal is to find a mapping function $T(X)$ that encodes point clouds into embeddings $\alpha$, such that Euclidean distances between encodings approximate the OT distances been point clouds. By computing the pairwise Wasserstein distance matrix for small mini-batches, we can train $T(X)$ so that encoding distances directly match Wasserstein.
Since $T(X)$ is applied to point clouds, it should contain a few key attributes:
  • Wasserstein distance is invariant with respect to the order of samples within point clouds ($W(X,PY)=W(X,Y)$, $P$ is a permutation matrix). The encoding must therefore also be permutation invariant: $T(X)=T(PX)$.
  • Point clouds contain varying numbers of samples. $T(X)$ must be able to generalize to arbitrary input length.
Transformers—specifically, stacked layers of multi-head attention—fit both of these criteria. Although they are commonly associated with sequential data such as words in a sentence, order information is provided as input to the model. Without the positional embeddings, attention mechanisms (and therefore transformers) are permutation equivariant (proof in supplementary section 8). By performing an invariant function such as averaging on the final set of embeddings, this produces a mapping $T(X)$ that is also invariant by design.
By training the encoder network $T(X)$ (Figure 1, Algorithm 1), we can produce encodings which allow for OT-based analysis of point clouds in linear time, enabling methods such as clustering and visualization [25] on the entire cohort. However, the encoder alone does not provide a clear interpretation of these analyses.
To imbue the embeddings with interpretability, we include a decoder network $G(\alpha)$ which takes encodings as input and is simultaneously trained to reproduce the input point clouds by minimizing the Wasserstein distance loss. Unlike standard autoencoders, which only train the full network to match outputs to inputs, Wormhole imposes a loss on both the encodings and decodings.
5 Embedding Non-Euclidean Distance Matrices
Given a distance matrix $D$ sized $N\times N$, we are interested in how well it can be approximated with a series of embeddings $\{\alpha_{i}\}_{i=1}^{N}$ in $d$ dimensions $\alpha_{i}\in R^{d}$. This corresponds to the optimization problem of minimizing the cumulative stress:
$\mathcal{L}=\min_{\{\alpha_{1},\alpha_{2},\dots,\alpha_{N}\}}\sum_{i,j}(\lVert\alpha_{i}-\alpha_{j}\rVert_{2}^{2}-D_{i,j})^{2}\quad$ (5)
Table 2: Comparison between embedding algorithms and theoretical bounds on small datasets where full Wasserstein distance matrices are realizable. Values represent the stress between the embedding and true distance matrices and parentheses denote the size of each dataset. Current MDS embedding algorithms like cMDS and non-linear SMACOF produce far from optimal errors, falling outside our bounds. Wormhole training is stochastic, yet it still converges between the error bounds and approaches the error obtained by the theoretically optimal PGD. Both PGD and Wormhole embeddings were well correlated ($\rho=0.999$) with OT distances for all datasets.
Dataset Lower Upper PGD (our) Wormhole (our) cMDS SMACOF
Simplex ($35$) $0.765$ $6.369$ $1.117$ $1.420$ $7.134$ $5.994$
Gaussian ($128$) $0.129$ $2.552$ $0.168$ $0.401$ $2.681$ $11.729$
MNIST ($256$) $0.042$ $0.616$ $0.058$ $0.100$ $0.657$ $3.231$
In the case where $D$ is Euclidean, by definition there exists a set of embeddings in dimension $d=N$ where the distance matrix is perfectly reconstructed and $\mathcal{L}=0$. Wormhole, however, is trained to minimize the stress relative to the Wasserstein distance matrix, which is generally not Euclidean. It is currently unclear to what extent $D$ can be represented with embeddings when it is not Euclidean.
As formulated in equation 5, this problem is non-convex and therefore difficult to study. However, when we instead reparameterize the objective to be a function over the distance matrix produced from the embedding $\hat{D}_{i,j}=\lVert\alpha_{i}-\alpha_{j}\rVert_{2}^{2}$ and add relevant constraints for $\hat{D}$ to be Euclidean, the problem is rendered convex:
$\begin{gathered}\mathcal{L}=\min_{\hat{D}}\sum_{i,j}(\hat{D}_{i,j}-D_{i,j})^{2}\\ \text{s.t. $C_{1}$: }-J\hat{D}J\text{ is PSD, $C_{2}$: }\hat{D}\text{ is hollow and $C_{3}$: }\hat{D}_{i,j}\geq 0.\end{gathered}$ (6)
[…]
Theorem 5.1 (Lower Bound).
For a given distance matrix $D$ and the eigendecomposition $\{\lambda_{i},v_{i}\}_{i=1}^{N}$ of its criterion matrix $F=-JDJ$, the optimal stress is greater or equal to:
[…]
This bound also appears in [15], which analyzed pathologies of cMDS on non-Euclidean distance matrices.
Theorem 5.2 (Upper Bound).
The optimal stress is less than or equal to:
[…]
Where:
$\Delta g_{i}=\frac{1}{2}\sum_{j:\lambda_{j}<0}\lambda_{j}\cdot v_{i,j}^{2}$ (9)
And $v_{i,j}$ is the $j$-th element of the $i$-th eigenvector of $F$.
The prominence of the negative spectrum of the criterion matrix can be thought of as a quantitative measure of how close $D$ is to being Euclidean. As we expect, the more minute the negative eigenvalues of $F$ (as in Figure S2), the smaller both the lower and upper bounds become, and we can anticipate that the embeddings will increase in quality.
Algorithm 2 Theoretically Optimal Embedding of non-Euclidean Distance Matrices with Projected Gradient Descent
  Input: Distance Matrix $D$, Initial Solution $D^{\prime}_{0}$, Learning Rate $\varepsilon$
  repeat
     $D^{*}_{t+1}=D^{\prime}_{t}-\varepsilon\nabla(\|D^{\prime}_{t}-D\|_{F}^{2})$GD Step}
     $D^{\prime}_{t+1}=\text{proj}_{C_{1}\cup C_{2}\cup C_{3}}(D^{*}_{t+1})$Dykstra}
  until Convergence
  RETURN $X=\text{cMDS}(D^{\prime})$Embed $D^{\prime}$ with MDS}
The derivation of the lower bound comes from removing the constraints $C_{1}$ and $C_{2}$ on hollowness and non-negativity of $\hat{D}$ and computing the projection operator onto $C_{1}$ which is the space of matrices where $-JDJ$ is positive semi-definite (subsection 9.1). Although it is similar to the standard projection onto PSD and incurs the same truncation error, it has additional terms due to the double-centering transform.
This new projection operator, along with the simple projections onto $C_{2}$ (hollow matrices, by subtraction of the diagonal) and $C_{3}$ (non-negative matrices, by thresholding), allows us to formulate a PGD [16, 40] scheme (Algorithm 2) to embed non-Euclidean distance matrices. This algorithm projects onto $C_{1}\cup C_{2}\cup C_{3}$ using Dykstra iterations [40] and is guaranteed to find the optimum for equation 6, unlike other MDS approaches such as cMDS or SMACOF [41]. The PGD algorithm requires knowing the full pairwise distance matrix and is therefore infeasible in our setting of massive cohorts. However, for small cohorts, the theoretical assurances of PGD provide a useful benchmark for Wormhole.
Our analysis indicates that in real-world datasets, the Wormhole approximation is highly accurate and commensurate with theoretical values. Although this assessment is based on a small subsample of the full dataset, we assume that if random subsets of the larger dataset can be well embedded, the embedding of the entire dataset is also reliable (Table 2). Full derivation of the theoretical bounds, and description of the PGD algorithm appear in Supplementary Sections 9 & 10.
Fig. 3: Wormhole on MNIST point clouds. a. Example of a point cloud produced by thresholding MNIST images. b. Retrieval of the Wasserstein distance in Wormhole embedding space. From a random sample of $128$ point clouds from a test set, we denote the correlation between the true pairwise OT distance and the Euclidean distance between their embeddings. The line $y=x$ is drawn in blue for reference. c. 2D UMAP visualization of Wormhole embeddings of the training and test set point clouds, which recapitulates ground-truth image digit without utilizing label information during training.
6 Experiments
We demonstrate Wasserstein Wormhole on datasets representing multiple contexts where OT is frequently applied. We applied Wormhole to 2D point clouds derived from the MNIST and Fashion-MNIST datasets, and to 3D CAD models from ModelNet40 and ShapeNet [42]. As a high-dimensional, real-world application, we used Wormhole to embed multicellular microenvironments [43] from two single-cell-resolution spatial transcriptomics datasets [44, 45], along with groups of cells (’MetaCells’ [46]) from a scRNA-seq atlas of patient response to COVID [47]).
Fig. 4: Generalization to OOD on Fashion-MNIST. During training, we held out out point clouds labeled ’Bag’. Performance was slightly lower compared to observed classes, especially for larger Wasserstein distances, as the MSE for OOD samples was $4.38\cdot 10^{-5}$ as opposed to $3.37\cdot 10^{-5}$ for observed samples. While not trained on any ’Bag’ point cloud, their encodings still largely agree with true OT distance, producing a Pearson correlation of $\rho=0.995$ and label accuracy of $0.97$.
For each dataset, we quantitatively benchmarked Wormhole by measuring how well the learned embeddings matched the true OT distance and ground-truth labels. Since the full Wasserstein distance matrix is unattainable, for each dataset we randomly sampled $10$ instances of $n=128$ point clouds and compared the $n\times n$ OT distance matrix to the embedding Euclidean distance. All the datasets are labeled, allowing us to measure the classification accuracy based on held-out test-sets for each.
We benchmarked our Wormhole against DiffusionEMD and Deep Wasserstein Embedding (DWE), both sharing Wormhole’s goal of finding OT preserving embeddings. We also compare against a constructed baseline approach we dub Fréchet [48] in which we use the closed-form $W_{2}$ distance for Gaussian distributions. For each point cloud, we calculated the mean ($\mu$) and covariance matrix ($\Sigma$) of its samples and use the following formula to approximate the OT distance between each pair of point clouds:
$W_{2}^{2}(X_{i},X_{j})\approx\|\mu_{i}-\mu_{j}\|^{2}+\text{Tr}\left(\Sigma_{i}+\Sigma_{j}-2\sqrt{\Sigma_{i}\Sigma_{j}}\right)$ (10)
The following sections present vignettes of Wormhole from several dataset, and results are summarized in Table 1. Wormhole outperforms all competing approaches and uniquely scales to massive & high-dimensional datasets such as the the 256,058-sample, 254-dimensional MERFISH Cell Niches cohort. We specifically highlight the improvement of Wormhole over DWE, which similarly learns a neural network parametric embedding. The transformer backbone of Wormhole proves to be superior to DWE’s CNN mode in low-dimensional dataset, while also essential in scaling Wormhole beyond 3D data.
6.1 MNIST (Touchstone)
We applied Wormhole to the MNIST dataset of hand-written digits, which consists of 70,000 samples of $28\times 28$ pixel grayscale images. We converted each sample to a point cloud by thresholding and selecting the set of pixel coordinates that passed the threshold (Figure 3a). While MNIST is considered small by modern machine-learning standards, straightforward computation of all $\binom{70000}{2}=2.45\cdot 10^{9}$ pairwise Wasserstein distances using current OT solvers is far from computationally feasible and would require a week of JAX-OTT [49] on an 80GB GPU (Figure 2).
By contrast, training Wormhole with default parameters (Supplementary Section 11) requires a small fraction of OT computations (approximately a $2000$-fold reduction) compared to the naive approach, with a total runtime of only 10 min. Despite the speedup, Wormhole still achieves remarkable performance, learning common patterns within the MNIST dataset and significantly outperforming contemporary approaches. Pairwise distances between point cloud encodings match their true OT distances and represent underlying digits (Figure 3b,c).
6.2 Fashion-MNIST (Out-Of-Distribution)
Similar to MNIST, FashionMNIST contains 70,000 images that each measure $28\times 28$ pixels, but the images depict articles of clothing instead of digits. After thresholding FashionMNIST images to produce point clouds, Wormhole was able to accurately recover OT distances and sample labels (Table 1) in the embedded space.
For a model to be deployed confidently, it is imperative that it can generalize to out-of-distribution (OOD) examples. To evaluate the capabilities of Wormhole on OOD data, we hid all examples labeled ’Bag’ during training and measured the quality of encodings. We found that Wormhole preserves the OT metric (Figure 4). In addition, when classifying test-set encodings based on in-distribution and OOD samples, accuracy does not diminish on the held-out class. Regardless of whether or not ’Bag’ samples were included in Wormhole training, their test-set accuracy was consistently $0.97$. We repeat this experiment for all $10$ classes in the FashionMNIST cohort and find that results are robust, having an overall OOD error of $3.14\cdot 10^{-}5\pm 2.22\cdot 10^{-5}$ and correlation of $0.998\pm 0.001$.
6.3 ModelNet40 (Variational Wasserstein)
The ModelNet40 dataset comprises point clouds from 3D synthetic CAD models of 40 diverse object classes. While each point cloud is more complex than the image datasets, containing thousands of points as opposed to hundreds, Wormhole was able to produce accurate embeddings in $30$ GPU min of training time (Table 1).
Fig. 5: Wormhole Variational Wasserstein. a. Mean encodings of each class were passed through the decoder to visualize their Wasserstein barycenters. Working in Wormhole space provides a quick and straightforward proxy for OT analysis, and recovered decodings reproduce the underlying class well. For brevity, we show 10 out of the 40 classes in from ModelNet40. b. Wormhole interpolation between two point clouds within the same class (top) and across different classes (bottom). Intermediate point clouds are linear interpolations between the embedding of each sample.
As a more challenging test case, we assessed the ability of the Wormhole decoder to interpret the embeddings and solve variational Wasserstein problems. The Wasserstein barycenters for a set of distributions is their average OT distance. Calculating barycenters is computationally challenging even in a small cohort. Wormhole embedding space is a Euclidean approximation of Wasserstein and barycenters of a set can be computed by calculating their encoding average and decoding it (Figure 5a).
Using Wormhole, we can also smoothly transition between point clouds via Wasserstein interpolation in the embedded space. Interpolating between distributions in Wasserstein space involves solving a weighted version of the barycenter problem as many times as there are intermediate samples. With Wormhole, we can leverage the Euclidean approximation of its latent space and directly decode linear interpolations between embeddings (Figure 5b). Despite the fact that interpolated embeddings are outside of the Wormhole training distribution, their decodings are valid and reliably visualize point cloud transitions.
6.4 ShapeNet (Gromov-
Wasserstein Wormhole)
Gromov-Wasserstein (GW) [50] is an extension of OT to distributions in different spaces that seeks the most isometric map between point clouds. Given two point clouds, their GW distance is:
$GW(X,Y)=\min_{P\in U(\nu,\mu)}\sum_{i,j,i^{\prime},j^{\prime}}\lVert d^{X}_{i,i^{\prime}}-d^{Y}_{j,j^{\prime}}\rVert^{2}P_{i,j}P_{i^{\prime},j^{\prime}}$ (11)
Where $d^{X}_{i,i^{\prime}}$ is the distance between $X_{i}$ and $X_{i^{\prime}}$, and similarly for $d^{Y}_{j,j^{\prime}}$. This problem is quadratic and thus, it is significantly more computationally intensive than standard OT [2]. GW has been a promising tool for aligning datasets from different modalities [51], mapping between graphs [52] and quantifying distances between 3D shapes with arbitrary orientation [53]. However, the utility of GW is curtailed by its complexity, motivating our Wormhole approach.
ShapeNet is a CAD-based dataset of 3D point clouds consisting of 16 different object classes. To construct the conditions where GW is usually applied, we randomly rotated each sample in 3D dimensions (Figure 6a). We then trained Wormhole to encode the rotated point clouds such that distances between embeddings retain GW distances.
Although the training objective was radically different, Wormhole generalized to GW and learned accurate embeddings (Table 1). In addition, the encodings were reflective of object classes despite input coordinates being randomly rotated, implying that GW Wormhole was able to produce rotation invariant embeddings (Figure 6b).
Fig. 6: Gromov-Wasserstein Wormhole. a. Examples of point clouds from the ShapeNet dataset randomly rotated in 3D to reflect common GW applications. b. Embeddings learned by GW Wormhole, visualized in 2D with UMAP and colored by object class. Despite their random orientations, Wormhole was able to encode the ShapeNet point clouds and represent their class, as well as recover GW distances (Table 1).
6.5 Cellular Niches (High-Dimensional Data)
Spatially resolved transcriptomics (ST) are a set of technologies which capture both RNA information and the spatial context of cells to gain insight into tissue organization. OT has been frequently applied [54, 55] to analyze results of ST and unearth the relationship between a cell’s phenotype and its microenvironment. For example, Spatial Optimal Transport (SPOT) [56] models the niche of each cell as the point cloud from the gene expression of every cell in its environment, and defines differences between niches using Wasserstein.
While OT is an alluring metric to quantify cellular niches, SPOT cannot scale to ST datasets of typical size. As experimental technologies mature, atlas-level spatial assays continue to grow in ubiquity and size. One instance is the $254$-gene MERFISH dataset of the motor cortex [44], which consists of 256,058 cells across 64 individual slices.
We summarize the $50\mu m$ radius around each cell in the MERFISH dataset to construct niche point clouds in $254$ dimensions. OT linearization methods such as DWE and DiffusionEMD assume that data is voxelized or discretized. While low-dimensional datasets can be manipulated to fit within their frameworks, the curse of dimensionality prevents adequate voxelization of this Cell Niches dataset. As commercial technologies capture more genes [57], the need for a method that can scale to higher dimensions, such as Wormhole, becomes more acute.
Parvalbumin-positive (Pvalb) neurons are one of four major groups of inhibitory neuron, and are crucial for the regulation of neural circuitry. They are sparse in the motor cortex, comprising only $3\%$ of the dataset. To understand their spatial organization, it is therefore critical to study the entire cohort and integrate insights from multiple slices.
Fig. 7: Wasserstein Wormhole on 254 dimensional spatial transcriptomics data. a. Visualization of Wormhole embeddings from niches of Pvalb neurons from the MERFISH dataset, colored by their first diffusion component. b. Spatial organization of Pvalb cells in a representative MERFISH slice. Wormhole embeddings reveal the cortical arrangement of these cells, as the first DC corresponds to neuronal depth.
After training Wormhole on the complete datasets (Table 1), we analyze the embeddings of the niches of Pvalb neurons. The first diffusion component (DC) [58] of Wormhole encodings aligns with their cortical depth across all slices within the dataset (Figure 7a-b). Despite the size and dimensionality of this dataset, Wormhole can still identify key biological signals across the cohort.
7 Discussion
We have introduced an approach, Wasserstein Wormhole, for embedding point clouds such that OT distances are preserved and can be approximated efficiently. Although OT is a popular approach, applying it to cohorts of distributions requires costly computation that scales quadratically with cohort size. Wormhole expands the functionality of OT to large datasets.
Encoding based on OT distances is theoretically challenged by the fact that Wasserstein is a non-Euclidean metric. To alleviate this concern, we derived theoretical lower and upper bounds that delimit how far non-Euclidean distance matrices can be from the embeddings. Our bounds grow smaller and tighter as the distance matrix approaches Euclidean. We further describe an algorithm based on projected gradient descent which is guaranteed to find an optimal embedding. On examples where the full Wasserstein distance matrix is tractable, Wormhole falls within the bounds and is close to the optimal error.
We applied our approach to six different settings, ranging from 2D point clouds to massive spatial transcriptomics datasets with hundreds of genes. Across all settings, Wormhole embeddings recapitulated the Wasserstein space, accurately encoded key aspects of the point clouds and outperformed competing approaches. Furthermore, Wormhole’s decoder recovered solutions to variational Wasserstein problems such as barycenters and point cloud interpolations. Wormhole seamlessly extends to other OT metrics, such as GW, and learns from arbitrary orientations.
Training Wasserstein Wormhole never exceeded 30 min on a single GPU, even on the 2048 points-per-sample ModelNet40 & ShapeNet datasets. Our algorithm is implemented in JAX [59] and integrates with OT tools (OTT-JAX) [49] for efficient Wasserstein distance calculations. Wormhole can be easily applied to any point cloud dataset via a pip installable package [^1] See tutorial in github.com/dpeerlab/WassersteinWormhole.
Several methods have been proposed to learn approximations for the Wasserstein distance, including those based on neural networks. Unlike prior methods, however, Wormhole makes no assumptions about empirical distributions and can be applied to general point clouds in arbitrary dimensions, while providing superior performance.
We envision several future directions for this work. The complexity of transformers is quadratic in input size, limiting their scalability to point clouds with more than a few thousand samples. Alternatively, recently proposed models [60] can efficiently learn from point clouds by leveraging local geometry. Fitting those models within the Wormhole framework could produce a more versatile approach.
Wasserstein is not only used to define a distance between distributions, but also to find the optimal mapping across them. An alluring Wormhole extension would be for it to produce the ideal matching in addition to estimating distance. This could possibly be achieved by exploiting the differentiability of Wormhole embeddings with respect to the point clouds.

### v4

@@
 By lending scalability and interpretability to OT approaches, Wasserstein Wormhole unlocks new avenues for data analysis in the fields of computational geometry and single-cell biology.
+Software is available at http://wassersteinwormhole.readthedocs.io/en/latest/.
+Machine Learning, ICML
 1 Introduction
 Many problems in statistical learning boil down to computing and comparing probability distributions that fit empirical data, and different approaches vary in their attention to distribution shape, spread, and degree of parameterization.
-Optimal transport methods attempt to find an ideal pairing between individual data points from two distributions [1] by representing distances as the minimum cost of mapping between (weighted) point clouds.
-While original OT solvers, which rely on linear programming, scale poorly [2], newer approaches can reduce the computational complexity and generalize OT to new domains such as computational biology [3, 4, 5], numerical geometry [6], chemistry [7], and image processing [8].
+Optimal transport (OT) methods attempt to find an ideal pairing between individual data points from two distributions (Villani et al., 2009) by representing distances as the minimum cost of mapping between (weighted) point clouds.
+While original OT solvers, which rely on linear programming, scale poorly (Peyré et al., 2019), newer approaches can reduce the computational complexity and generalize OT to new domains such as computational biology (Nitzan et al., 2019; Schiebinger et al., 2019; Bunne et al., 2023), numerical geometry (Su et al., 2015), chemistry (Wu et al., 2023), and image processing (Feydy et al., 2017).
 We are specifically interested in massive cohorts of empirical distributions, each with up to thousands of points.
-The computational complexity of finding the Wasserstein distance between two point clouds with Sinkhorn’s algorithm [9], the current favored [10] OT solver, is $O(n^{2}\cdot I)$, where $n$ is the size of each point cloud and $I$ is the number of Sinkhorn iterations until convergence.
-Despite improvements in scalability, current OT algorithms are unable to analyze large cohorts as we must repeat Sinkhorn’s algorithm for all pairs in the cohorts, requiring $O(N^{2}\cdot n^{2}\cdot I)$ time to analyze $N$ distributions.
+The computational complexity of finding the Wasserstein distance between two point clouds with Sinkhorn’s algorithm (Cuturi, 2013), the current favored (Flamary et al., 2021) OT solver, is $O(n^{2}\cdot I)$, where $n$ is the size of each point cloud and $I$ is the number of Sinkhorn iterations until convergence.
+As we must repeat Sinkhorn’s algorithm for all pairs in a cohort, requiring $O(N^{2}\cdot n^{2}\cdot I)$ time to analyze $N$ distributions, current OT algorithms are unable to analyze large cohorts despite improvements in scalability.
 Here, we present Wasserstein Wormhole, an algorithm that represents each point cloud as a single embedded point, such that the Euclidean distance in the embedding space matches the OT distance between point clouds.
@@
 Point clouds contain a variable number of unordered samples, hence neural architectures that require a fixed-sized or ordered input are not suited to encode them.
-On the contrary, transformers [11] are uniquely poised to do so, due to their permutation equivariance, indifference to input length, and ability to model relationships between samples in each point cloud via attention.
-In addition to an encoder that embeds point clouds, Wormhole trains a decoder that reconstructs point clouds from their embedding.
-The decoder facilitates interpretability with its ability to explicitly compute Wasserstein barycenters [12] or interpolate between point clouds [13].
+On the contrary, transformers (Vaswani et al., 2017) are uniquely poised to do so, due to their permutation equivariance, indifference to input length, and ability to model relationships between samples in each point cloud via attention.
 Fig.
@@
 The Wormhole decoder (not shown) is a second transformer trained to reproduce the input point clouds from the embedding by minimizing the OT distance between input and output.
-The problem solved by Wormhole is analogous to multidimensional scaling [14].
+In addition to an encoder that embeds point clouds, Wormhole trains a decoder that reconstructs point clouds from their embedding.
+The decoder facilitates interpretability with its ability to explicitly compute Wasserstein barycenters (Cuturi & Doucet, 2014) or interpolate between point clouds (Chewi et al., 2021).
+The problem solved by Wormhole is analogous to multidimensional scaling (Torgerson, 1952).
 Both methods find an embedding that preserves distances between samples.
-While several distance metrics are provably capable of being represented in this manner, it has been shown that a host of other ’non-Euclidean’ metrics, including Wasserstein, cannot be exactly equated with Euclidean distance.
-Counter-intuitively, classical MDS on non-Euclidean distance matrices increases in error as embedding dimensions increase [15].
-Inspired by [15], we extend MDS theory to derive upper and lower bounds on the error incurred by our gradient descent-based approach, thereby providing Wormhole with theoretical guarantees on performance for (non-Euclidean) Wasserstein distances.
-We further describe a projected gradient descent (PGD) [16] algorithm for MDS which is guaranteed to converge to the global optimum for any distance matrix.
-On small examples where it is feasible to compute bounds and apply PGD, we empirically show that Wormhole approaches the best solution.
+While several distance metrics are provably capable of being represented in this manner, it has been shown that a host of other ”non-Euclidean” metrics, including Wasserstein, cannot be exactly equated with Euclidean distance.
+Counter-intuitively, classical MDS on non-Euclidean distance matrices becomes more inaccurate as embedding dimensions increase (Sonthalia et al., 2021).
+Inspired by (Sonthalia et al., 2021), we extend MDS theory to derive upper and lower bounds for the error incurred by embedding non-Euclidean distances, such as Wasserstein.
+We further describe a projected gradient descent (PGD) (Boyd & Vandenberghe, 2004) algorithm for MDS which is guaranteed to converge to the global optimum for any distance matrix.
+On small examples where it is feasible to compute bounds and apply PGD, we find that Wormhole approaches the optimal solution and falls within our derived bounds.
+Dataset Parameters OT Correlation Label Accuracy
+Name Cohort Size Cloud Size Dimension Wormhole DiffusionEMD DWE Fréchet Wormhole DiffusionEMD DWE Fréchet
+MNIST 70,000 105 2 0.98 0.845 0.92 0.86 0.98 0.84 0.97 0.49
+Fashion-MNIST 70,000 356 2 0.99 0.97 0.98 0.99 0.87 0.77 0.87 0.73
+ModelNet40 12,311 2048 3 0.99 0.85 0.97 0.95 0.86 0.69 0.78 0.61
+ShapeNet 15,011 2048 3 0.99 0.81 0.97 0.97 0.98 0.94 0.97 0.92
+Rotated ShapeNet (GW) 15,011 512 3 0.98 NA NA 0.68 0.82 NA NA 0.16
+MERFISH Cell Niches 256,058 11 254 0.97 OOM OOM OOM 0.98 OOM OOM OOM
+SeqFISH Cell Niches 57,407 29 351 0.98 OOM OOM OOM 0.97 OOM OOM OOM
+scRNA-seq Atlas 2,185 69 2500 0.98 OOM OOM OOM 0.96 OOM OOM OOM
+Table 1: Summary of each dataset used in this study and results from every benchmarked algorithm.
+Cloud size denotes the median number of points in each point cloud for every dataset.
+Pearson correlations are computed between true and estimated Wasserstein distances according to pairwise matrices of $10$ instances of $128$ randomly chosen point clouds from each dataset (OT Correlation).
+The mean squared error (MSE) follows the same trend with Wormhole outperforming competing methods (Table 3).
+Object classification accuracy is based on the embeddings of each algorithm.
+After computing encodings, a classifier was trained on the training set embeddings and we denote the classifier’s accuracy on the test set (Label Accuracy).
+Rotated ShapeNet is a version of the standard ShapeNet dataset where each sample is randomly rotated in 3D, subsampled by a factor of $4$, and embedded via Gromov-Wasserstein (GW).
+We divide the results into Wasserstein on 2D and 3D point-clouds, Gromov-Wasserstein on rotated data with arbitrary orientation and high-dimensional datasets.
+Full implementation details are described in the supplement (Section 11).
+NA, Not Applicable.
+OOM, Out-Of-Memory.
 2 Prior Work
-Learning from point-cloud data is dynamic area of research within machine learning [17], with many Optimal Transport based application.
-PointNet [18, 19] introduced the idea of permutation-invariant fully-connected neural networks to classify and segment 3D point-clouds, which was later extended into transformer based networks [20].
-In the context of OT, several works have noted the expressiveness of Wasserstein space compared to euclidean space [21, 22, 23], and have developed methods to embed point-clouds into Wasserstein space, from which relevant information can be more reliably extracted.
-These works are in essence the inverse of Wormhole, as their target space is Wasserstein, where in Wormhole, target space is euclidean.
-Wasserstein Wormhole belongs to a class of algorithms that exploit structures within distributions or relationships between cohorts of them to accelerate or approximate OT.
-Assuming the transport plan between two distribution is structured, Low-Rank (LR) Sinkhorn [24] factorizes the transport matrix, forcing it into a low-rank structure which enables faster computation of Sinkhorn iterations.
-Similarly, Convolutional and Geodesic Sinkhorn [25, 26] construct solvers which leverage the underlying geometry of the point-cloud space to simplify OT optimization.
-MetaOT [27] recognizes that convergence of Sinkhorn depends greatly on quality of initialization and learns a neural network which learns an initial Sinkhorn potential for a given pair of input distributions.
-Due to its fully-connected architecture, MetaOT is constrained to distributions which share supports, and can not be applied to arbitrary point-clouds.
-Although these methods can speed up the computation of an individual OT problem, none alleviate the need for many expensive pairwise computations, and are far slower than Wormhole on large cohorts, as shown in Figure 2.
-Like Wormhole, methods such as Deep Wasserstein Embeddings (DWE) [28] and DiffusionEMD [29] produce embeddings that preserve Wasserstein distances.
-DWE learns a neural network via a Siamese architecture trained to match distances, and DiffusionEMD extends the wavelet-based approximation of EMD [30] onto graph domains.
-Unlike our Wormhole approach, which applies to point clouds in any dimension, DWE is constrained to small 2D images or 3D grids, while scalability of DiffusionEMD is limited to distributions that share supports, such as grayscale images or voxelized spaces.
-We benchmark Wormhole against DiffusionEMD & DWE, and show that our method is superior at matching OT distances and deriving insights from embeddings, and Wormhole is the only embedding method which scales to high-dimensional datasets (Table 1).
+There have been many prior works on learning from point cloud data (Guo et al., 2020), often involving OT-based solutions.
+PointNet (Garcia-Garcia et al., 2016; Qi et al., 2017) introduced the idea of permutation-invariant, fully-connected neural networks to classify and segment 3D point-clouds, which was later extended into transformer-based networks (Lee et al., 2019).
+In the context of OT, several works have noted the expressiveness of Wasserstein space compared to euclidean space (Frogner et al., 2019; Katageri et al., 2024; Liu et al., 2022), and have developed methods to embed point-clouds into Wasserstein space, from which relevant information can be more reliably extracted.
+These works are in essence the inverse of Wormhole, as their target space is Wasserstein, in contrast to Wormhole’s Euclidean target space.
 Fig.
@@
 However, even in cohorts with relatively few samples, Wormhole is superior and no other method can scale to complete datasets, requiring weeks of compute time, even on a fully-utilized 80GB GPU.
-Wormhole follows a long line of neural network approaches for Optimal Transport estimation [31].
-Sinkhorn iterations, which are quadratic with size of point-clouds, cannot be applied when When point-clouds are massive consisting of thousands of points.
-Neural OT [32] parameterizes the transport plan between two empirical distributions as a input-convex neural network [33] which allows for fast, mini-batch training.
-Such models have been incredibly successful at extending OT to scRNA-seq data [34, 5, 35], where samples commonly contain over $10^{5}$ cells.
-Wormhole works best in a different setting, where individual OT problems can be solved, but their solvers are not efficient enough to scale to the many pairwise comparisons within a large cohort.
-Future extensions of our work can make use of neural solvers to unlock Wasserstein analysis to large cohorts of massive point-clouds.
+Wasserstein Wormhole belongs to a class of algorithms that exploit structures within distributions or relationships between cohorts of them to accelerate or approximate OT.
+Low-Rank (LR) Sinkhorn (Scetbon et al., 2021) factorizes the pairing matrix into a low-rank composition, assuming the transport plan between two distribution is structured, to enable faster computation of Sinkhorn iterations.
+Convolutional and Geodesic Sinkhorn (Bachmann et al., 2022; Solomon et al., 2015) construct solvers which leverage the underlying geometry of the point-cloud space to simplify OT optimization.
+MetaOT (Amos et al., 2022) recognizes that convergence of Sinkhorn depends greatly on quality of initialization and fits a neural network which learns an initial Sinkhorn potential for a given pair of input distributions.
+Due to its fully-connected architecture, MetaOT is constrained to distributions which share supports, and cannot be applied to arbitrary point-clouds.
+Although these methods can speed up the computation of an individual OT problem, none alleviate the need for many expensive pairwise computations, and are far slower than Wormhole on large cohorts (Figure 2).
+Previous works have also employed Wasserstein preserving euclidean embeddings in the pursuit of scaling OT to large cohorts (Tong et al., 2021), some using neural network models (Courty et al., 2017; Sehanobish et al., 2020).
+DWE (Courty et al., 2017) trains a Siamese neural network to match distances, and DiffusionEMD (Tong et al., 2021) extends the wavelet-based approximation of EMD (Shirdhonkar & Jacobs, 2008) onto graph domains.
+Unlike our Wormhole approach, which applies to point clouds in any dimension, DWE is constrained to small 2D images or 3D grids, while scalability of DiffusionEMD is limited to distributions that share supports, such as grayscale images or voxelized spaces.
+We benchmark Wormhole against DiffusionEMD & DWE, and show that our method is superior at matching OT distances and deriving insights from embeddings on low-dimensional point clouds.
+On general datasets, Wormhole is the only embedding method which can scale to high-dimensions.
 3 Preliminaries
 3.1 Optimal Transport
-We follow the formulation of optimal transport presented in [2].
+We follow the formulation of optimal transport presented in (Peyré et al., 2019).
 Given two discrete distributions $P_{x}=\sum_{i=1}^{n}\mu_{i}\delta_{x_{i}}$ and $P_{y}=\sum_{i=1}^{m}\nu_{i}\delta_{y_{i}}$ ($\delta$ is the Dirac function), their entropically regularized OT distance is:
@@
 Where $U(\nu,\mu):=\{P\in R_{+}^{n,m}:\sum_{j}P_{i,j}=\nu_{i},\sum_{i}P_{i,j}=\mu_{j}\}$ is the set of all valid pairing matrices, $c(x_{i},y_{j})$ is the ground cost function between points in $P_{x}$ and $P_{y}$, usually chosen to be the Euclidean distance, and $H(P)=-\sum_{i,j}P_{i,j}(\log P_{i,j}-1)$ is the entropic term.
-While the unregularized problem incurs a heavy computational burden for distributions with large supports or in high dimensions, we can compute $W_{\varepsilon}$ via a series of GPU-accelerable vector-matrix products known as Sinkhorn iterations [9].
+While the unregularized problem incurs a heavy computational burden for distributions with large supports or in high dimensions, $W_{\varepsilon}$ can be computed via a series of GPU-accelerable vector-matrix products known as Sinkhorn iterations (Cuturi, 2013).
 Due to the regularization term, the $W_{\varepsilon}$ distance between a distribution and itself is not 0, which can introduce biases when used to optimize generative models.
-To allow for effective optimization while retaining the computational advantages of entropic regularization, a common alternative is the Sinkhorn Divergence [36], which subtracts the self-transport costs from $W_{\varepsilon}$:
+To allow for effective optimization while retaining the computational advantages of entropic regularization, a common alternative is the Sinkhorn Divergence (Genevay et al., 2018), which subtracts the self-transport costs from $W_{\varepsilon}$:
 $S_{\varepsilon}(P_{x},P_{y})=W_{\varepsilon}(P_{x},P_{y})-\frac{W_{\varepsilon}(P_{x},P_{x})+W_{\varepsilon}(P_{y},P_{y})}{2}$ (2)
@@
 $\forall{i,j}:\;D_{i,j}=\|x_{i}-x_{j}\|_{2}^{2}$ (3)
-From [37], a matrix $D$ is Euclidean if and only if all of its values are non-negative, it is hollow (diagonals are 0) and the criterion matrix $F$ is positive-semidefinite (PSD):
+From (Gower, 1985), a matrix $D$ is a Euclidean Distance Matrix if and only if all of its values are non-negative, it is hollow (diagonal elements are 0) and the criterion matrix $F$ is positive-semidefinite (PSD):
 $F=-JDJ$ (4)
 Where $J=I_{N}-\frac{\mathbf{1}_{N}\mathbf{1}_{N}^{\top}}{N}$ is the centering matrix and $1_{N}$ is a vector of length $N$ filled with ones.
-From brevity, we will call matrices which fit these criterion Euclidean.
-Wasserstein distance is not Euclidean (see section 8.3 in [2] and [38]), and there are many examples of cohorts of distributions (Figures S1,S2) wherein pairwise OT distance matrices do not fit the criteria in equation 4.
-Unlike the Euclidean scenario where classical MDS (cMDS) improves as embedding size grows, on non-Euclidean distance matrices the error behaves counter-intuitively and worsens as dimension increases [15].
+Wasserstein distance is not Euclidean (see section 8.3 in (Peyré et al., 2019) and (Naor & Schechtman, 2007)), and there are many examples of cohorts of distributions (Figures S1,S2) wherein pairwise OT distance matrices do not fit the criteria in equation 4.
+Unlike the Euclidean scenario where classical MDS (cMDS) improves as embedding size grows, on non-Euclidean distance matrices the error behaves counter-intuitively and worsens as dimension increases (Sonthalia et al., 2021).
 This motivates us to seek a description of the theory underpinning the embeddings of non-Euclidean distances.
-Dataset Parameters OT Correlation Label Accuracy
-Name Cohort Size Cloud Size Dimension Wormhole DiffusionEMD DWE Fréchet Wormhole DiffusionEMD DWE Fréchet
-MNIST 70,000 105 2 0.986 0.845 0.923 0.862 0.98 0.84 0.97 0.49
-Fashion-MNIST 70,000 356 2 0.999 0.971 0.985 0.992 0.87 0.77 0.87 0.73
-ModelNet40 12,311 2048 3 0.996 0.852 0.979 0.952 0.86 0.69 0.78 0.61
-ShapeNet 15,011 2048 3 0.995 0.817 0.977 0.973 0.98 0.94 0.97 0.92
-Rotated ShapeNet (GW) 15,011 512 3 0.987 NA NA 0.687 0.82 NA NA 0.16
-MERFISH Cell Niches 256,058 11 254 0.977 OOM OOM OOM 0.98 OOM OOM OOM
-SeqFISH Cell Niches 57,407 29 351 0.984 OOM OOM OOM 0.97 OOM OOM OOM
-scRNA-seq Atlas 2,185 69 2500 0.986 OOM OOM OOM 0.96 OOM OOM OOM
-Table 1: Summary of each dataset used in this study and results from every benchmarked algorithm.
-Cloud size denotes the median number of points in each point cloud for every dataset.
-Pearson correlations are computed between true and estimated Wasserstein distances according to pairwise matrices of $10$ instances of $128$ randomly chosen point clouds from each dataset (OT Correlation).
-The mean squared error (MSE) follows the same trend with Wormhole outperforming competing methods (Table 3).
-Object classification accuracy is based on the embeddings of each algorithm.
-After computing encodings, a classifier was trained on the training set embeddings and we denote the classifier’s accuracy on the test set (Label Accuracy).
-Rotated ShapeNet is the ShapeNet dataset where each sample is randomly rotated in 3D, subsampled by a factor of $4$, and embedded via Gromov-Wasserstein (GW).
-We divide the results into Wasserstein on 2D and 3D point-clouds, Gromov-Wasserstein on rotated data with arbitrary orientation and high-dimensional datasets.
-Full implementation details are described in the supplement (Section 11).
-NA, Not Applicable.
-OOM, Out-Of-Memory.
 4 Wasserstein Wormhole
+The complexity of modern OT solvers is quadratic with respect to the number of points in a point cloud; thus, it is only straightforward to compute pairwise Wasserstein distance for small cohorts of empirical distributions.
+Indeed, naive approaches will not even scale to standard datasets, such as the 12,311-point-cloud ModelNet40 benchmark (Wu et al., 2015).
+This problem calls for a stochastic approach.
+Instead of directly assigning each point cloud with an embedding, we propose learning a parametric encoding function.
+Formally, the dataset in our setting is a cohort of point clouds $\Omega=\{X_{1},X_{2},X_{3},\dots X_{N}\}$, where each point cloud $X_{i}$ contains $n_{i}$ points in $d$ dimensions.
+Our goal is to find a mapping function $T(X)$ that encodes point clouds into embeddings $\alpha$, such that Euclidean distances between encodings approximate the OT distances been point clouds.
+By computing the pairwise Wasserstein distance matrix for small mini-batches, we can train $T(X)$ so that encoding distances directly match Wasserstein.
+Since $T(X)$ is applied to point clouds, it should contain a few key attributes:
+  • Wasserstein distance is invariant with respect to the order of samples within point clouds ($W(X,PY)=W(X,Y)$, $P$ is a permutation matrix).
+The encoding must therefore also be permutation invariant: $T(X)=T(PX)$.
+  • Point clouds contain varying numbers of samples. $T(X)$ must be able to generalize to arbitrary cloud sizes while modeling relationships between the points within them.
+Transformers—specifically, stacked layers of multi-head attention—fit both of these criteria.
+Although they are commonly associated with sequential data such as words in a sentence, order information is provided as input to the model.
+Without the positional embeddings, attention mechanisms (and therefore transformers) are permutation equivariant (proof in supplementary section 8).
+By performing an invariant function such as averaging on the final set of embeddings, this produces a mapping $T(X)$ that is also invariant by design.
 Algorithm 1 Wasserstein Wormhole At each training step, a mini-batch of point clouds is randomly sampled from the cohort.
 Wormhole is optimized so that pairwise distance between mini-batch embeddings match their OT distance and output decodings reconstruct input point clouds.
-  Input: point clouds $\Omega=\{X_{i}\}^{N}_{i=1}$, initialized encoder $T$ and decoder $G$, batch size $B$, learning rate $\varepsilon_{t}$
+  Input: point clouds $\Omega=\{X_{i}\}^{N}_{i=1}$, initialized encoder $T$ and decoder $G$ networks, batch size $B$, learning rate $\varepsilon_{t}$
   repeat
@@
   until Convergence
-The complexity of modern OT solvers is quadratic with respect to the number of points in a point cloud; thus, it is only straightfoward to compute pairwise Wasserstein distance for small cohorts of empirical distributions.
-Indeed, naive approaches will not even scale to standard datasets, such as the 12,311-point-cloud ModelNet40 benchmark [39].
-This problem calls for a stochastic approach.
-Instead of directly assigning each point cloud with an embedding, we propose learning a parametric encoding function.
-Formally, the dataset in our setting is a cohort of point clouds $\Omega=\{X_{1},X_{2},X_{3},\dots X_{N}\}$, where each point cloud $X_{i}$ contains $n_{i}$ points in $d$ dimensions.
-Our goal is to find a mapping function $T(X)$ that encodes point clouds into embeddings $\alpha$, such that Euclidean distances between encodings approximate the OT distances been point clouds.
-By computing the pairwise Wasserstein distance matrix for small mini-batches, we can train $T(X)$ so that encoding distances directly match Wasserstein.
-Since $T(X)$ is applied to point clouds, it should contain a few key attributes:
-  • Wasserstein distance is invariant with respect to the order of samples within point clouds ($W(X,PY)=W(X,Y)$, $P$ is a permutation matrix).
-The encoding must therefore also be permutation invariant: $T(X)=T(PX)$.
-  • Point clouds contain varying numbers of samples. $T(X)$ must be able to generalize to arbitrary input length.
-Transformers—specifically, stacked layers of multi-head attention—fit both of these criteria.
-Although they are commonly associated with sequential data such as words in a sentence, order information is provided as input to the model.
-Without the positional embeddings, attention mechanisms (and therefore transformers) are permutation equivariant (proof in supplementary section 8).
-By performing an invariant function such as averaging on the final set of embeddings, this produces a mapping $T(X)$ that is also invariant by design.
-By training the encoder network $T(X)$ (Figure 1, Algorithm 1), we can produce encodings which allow for OT-based analysis of point clouds in linear time, enabling methods such as clustering and visualization [25] on the entire cohort.
+By training the encoder network $T(X)$ (Figure 1, Algorithm 1), we can produce encodings which allow for OT-based analysis of point clouds in linear time, enabling methods such as clustering and visualization (Bachmann et al., 2022) on the entire cohort.
 However, the encoder alone does not provide a clear interpretation of these analyses.
+Fig.
+3: Wormhole on MNIST point clouds.
+a.
+Example of a point cloud produced by thresholding MNIST images.
+b.
+Retrieval of the Wasserstein distance in Wormhole embedding space.
+From a random sample of $128$ point clouds from the test set, we denote the correlation between the true pairwise OT distance and the Euclidean distance between their embeddings.
+The line $y=x$ is drawn in blue for reference.
+c.
+2D UMAP visualization of Wormhole embeddings of the training and test set point clouds, which recapitulates ground-truth image digit without utilizing label information during training.
 To imbue the embeddings with interpretability, we include a decoder network $G(\alpha)$ which takes encodings as input and is simultaneously trained to reproduce the input point clouds by minimizing the Wasserstein distance loss.
@@
 $\mathcal{L}=\min_{\{\alpha_{1},\alpha_{2},\dots,\alpha_{N}\}}\sum_{i,j}(\lVert\alpha_{i}-\alpha_{j}\rVert_{2}^{2}-D_{i,j})^{2}\quad$ (5)
-Table 2: Comparison between embedding algorithms and theoretical bounds on small datasets where full Wasserstein distance matrices are realizable.
-Values represent the stress between the embedding and true distance matrices and parentheses denote the size of each dataset.
-Current MDS embedding algorithms like cMDS and non-linear SMACOF produce far from optimal errors, falling outside our bounds.
-Wormhole training is stochastic, yet it still converges between the error bounds and approaches the error obtained by the theoretically optimal PGD.
-Both PGD and Wormhole embeddings were well correlated ($\rho=0.999$) with OT distances for all datasets.
-Dataset Lower Upper PGD (our) Wormhole (our) cMDS SMACOF
-Simplex ($35$) $0.765$ $6.369$ $1.117$ $1.420$ $7.134$ $5.994$
-Gaussian ($128$) $0.129$ $2.552$ $0.168$ $0.401$ $2.681$ $11.729$
-MNIST ($256$) $0.042$ $0.616$ $0.058$ $0.100$ $0.657$ $3.231$
 In the case where $D$ is Euclidean, by definition there exists a set of embeddings in dimension $d=N$ where the distance matrix is perfectly reconstructed and $\mathcal{L}=0$.
 Wormhole, however, is trained to minimize the stress relative to the Wasserstein distance matrix, which is generally not Euclidean.
-It is currently unclear to what extent $D$ can be represented with embeddings when it is not Euclidean.
+It is currently unclear to what extent $D$ can be represented with embeddings when it is not Euclidean (Sonthalia et al., 2021).
 As formulated in equation 5, this problem is non-convex and therefore difficult to study.
@@
 […]
-This bound also appears in [15], which analyzed pathologies of cMDS on non-Euclidean distance matrices.
+This bound also appears in (Sonthalia et al., 2021), which analyzed pathologies of cMDS on non-Euclidean distance matrices.
 Theorem 5.2 (Upper Bound).
@@
 As we expect, the more minute the negative eigenvalues of $F$ (as in Figure S2), the smaller both the lower and upper bounds become, and we can anticipate that the embeddings will increase in quality.
-Algorithm 2 Theoretically Optimal Embedding of non-Euclidean Distance Matrices with Projected Gradient Descent
-  Input: Distance Matrix $D$, Initial Solution $D^{\prime}_{0}$, Learning Rate $\varepsilon$
+The derivation of the lower bound comes from removing the constraints $C_{1}$ and $C_{2}$ on hollowness and non-negativity of $\hat{D}$ and computing the projection operator onto $C_{1}$ which is the space of matrices where $-JDJ$ is positive semi-definite (subsection 9.1).
+Although it is similar to the standard projection onto PSD and incurs the same truncation error, it has additional terms due to the double-centering transform.
+Algorithm 2 Theoretically Optimal Embedding of non-Euclidean Distance Matrices with Projected Gradient Descent (PGD)
+  Input: Distance Matrix $D$, Initial Solution $D^{\prime}_{0}$, Learning Rate $\varepsilon_{t}$
   repeat
-     $D^{*}_{t+1}=D^{\prime}_{t}-\varepsilon\nabla(\|D^{\prime}_{t}-D\|_{F}^{2})$GD Step}
+     $D^{*}_{t+1}=D^{\prime}_{t}-\varepsilon_{t}\nabla(\|D^{\prime}_{t}-D\|_{F}^{2})$GD Step}
      $D^{\prime}_{t+1}=\text{proj}_{C_{1}\cup C_{2}\cup C_{3}}(D^{*}_{t+1})$Dykstra}
@@
   RETURN $X=\text{cMDS}(D^{\prime})$Embed $D^{\prime}$ with MDS}
-The derivation of the lower bound comes from removing the constraints $C_{1}$ and $C_{2}$ on hollowness and non-negativity of $\hat{D}$ and computing the projection operator onto $C_{1}$ which is the space of matrices where $-JDJ$ is positive semi-definite (subsection 9.1).
-Although it is similar to the standard projection onto PSD and incurs the same truncation error, it has additional terms due to the double-centering transform.
-This new projection operator, along with the simple projections onto $C_{2}$ (hollow matrices, by subtraction of the diagonal) and $C_{3}$ (non-negative matrices, by thresholding), allows us to formulate a PGD [16, 40] scheme (Algorithm 2) to embed non-Euclidean distance matrices.
-This algorithm projects onto $C_{1}\cup C_{2}\cup C_{3}$ using Dykstra iterations [40] and is guaranteed to find the optimum for equation 6, unlike other MDS approaches such as cMDS or SMACOF [41].
+This new projection operator, along with the simple projections onto $C_{2}$ (hollow matrices, by setting the diagonal to 00) and $C_{3}$ (non-negative matrices, by thresholding), allows us to formulate a PGD (Boyd & Vandenberghe, 2004) scheme (Algorithm 2) to embed non-Euclidean distance matrices.
+This algorithm projects onto $C_{1}\cup C_{2}\cup C_{3}$ using Dykstra iterations (Boyle & Dykstra, 1986) and is guaranteed to find the optimum for equation 6, unlike other MDS approaches such as cMDS or SMACOF (De Leeuw & Mair, 2009).
 The PGD algorithm requires knowing the full pairwise distance matrix and is therefore infeasible in our setting of massive cohorts.
@@
 Our analysis indicates that in real-world datasets, the Wormhole approximation is highly accurate and commensurate with theoretical values.
-Although this assessment is based on a small subsample of the full dataset, we assume that if random subsets of the larger dataset can be well embedded, the embedding of the entire dataset is also reliable (Table 2).
-Full derivation of the theoretical bounds, and description of the PGD algorithm appear in Supplementary Sections 9 & 10.
-Fig.
-3: Wormhole on MNIST point clouds.
-a.
-Example of a point cloud produced by thresholding MNIST images.
-b.
-Retrieval of the Wasserstein distance in Wormhole embedding space.
-From a random sample of $128$ point clouds from a test set, we denote the correlation between the true pairwise OT distance and the Euclidean distance between their embeddings.
-The line $y=x$ is drawn in blue for reference.
-c.
-2D UMAP visualization of Wormhole embeddings of the training and test set point clouds, which recapitulates ground-truth image digit without utilizing label information during training.
+Although this assessment is based on a small subsample of the full dataset, we assume that if random subsets of the larger dataset can be well embedded, the embedding of the entire dataset is reliable (Table 2).
+Full derivation of the theoretical bounds appear in Supplementary Sections 9 & 10.
+Dataset Lower Upper PGD Wormhole cMDS
+Simplex ($35$) $0.765$ $6.369$ $1.117$ $1.420$ $7.134$
+Gaussian ($128$) $0.129$ $2.552$ $0.168$ $0.401$ $2.681$
+MNIST ($256$) $0.042$ $0.616$ $0.058$ $0.100$ $0.657$
+Table 2: Comparison between embedding algorithms and theoretical bounds on small datasets where full Wasserstein distance matrices are realizable.
+Values are the stress between the embedding and true distances and parentheses denote the size of each dataset.
+Classical MDS produces far from optimal errors, falling outside our bounds.
+Wormhole training is stochastic, yet it still converges between the error bounds and approaches the error obtained by the theoretically optimal PGD.
+Both PGD and Wormhole embeddings were well correlated ($\rho=0.99$) with OT distances.
 6 Experiments
 We demonstrate Wasserstein Wormhole on datasets representing multiple contexts where OT is frequently applied.
-We applied Wormhole to 2D point clouds derived from the MNIST and Fashion-MNIST datasets, and to 3D CAD models from ModelNet40 and ShapeNet [42].
-As a high-dimensional, real-world application, we used Wormhole to embed multicellular microenvironments [43] from two single-cell-resolution spatial transcriptomics datasets [44, 45], along with groups of cells (’MetaCells’ [46]) from a scRNA-seq atlas of patient response to COVID [47]).
+We applied Wormhole to 2D point clouds derived from the MNIST and Fashion-MNIST datasets, and to 3D CAD models from ModelNet40 and ShapeNet (Chang et al., 2015).
+As a high-dimensional applications, we used Wormhole to embed multicellular niches (Haviv et al., 2023) from two high-resolution spatial transcriptomics datasets (Zhang et al., 2021; Lohoff et al., 2022), along with groups of cells (’MetaCells’ (Persad et al., 2023)) from a scRNA-seq atlas of patient response to COVID (Stephenson et al., 2021).
+For each dataset, we quantitatively benchmarked Wormhole by measuring how well the learned embeddings matched the true OT distance and ground-truth labels.
+Since the full Wasserstein distance matrix is unattainable, for each dataset we randomly sampled $10$ instances of $n=128$ point clouds and compared the $n\times n$ OT distance matrix to the embedding Euclidean distance.
+All the datasets are labeled, allowing us to measure the classification accuracy based on held-out test-sets for each.
+We benchmarked our Wormhole against DiffusionEMD and Deep Wasserstein Embedding (DWE), both sharing Wormhole’s goal of finding OT preserving embeddings.
+We also compare against a constructed baseline we dub Fréchet (Dowson & Landau, 1982) in which we use the closed-form $W_{2}$ distance for Gaussian distributions.
+For each point cloud, we calculated the mean ($\mu$) and covariance matrix ($\Sigma$) of its samples and use the following formula to approximate the OT distance between each point cloud pair:
+$W_{2}^{2}\approx\|\mu_{i}-\mu_{j}\|^{2}+\text{Tr}\left(\Sigma_{i}+\Sigma_{j}-2\sqrt{\Sigma_{i}\Sigma_{j}}\right)$ (10)
+The following sections present vignettes of Wormhole from several dataset, and results are summarized in Table 1.
+Wormhole outperforms all competing approaches and uniquely scales to massive & high-dimensional datasets such as the the 256,058-sample, 254-dimensional MERFISH Cell Niches cohort.
+We specifically highlight the improvement of Wormhole over DWE, which similarly learns a neural network parametric embedding.
+The transformer backbone of Wormhole is superior to DWE’s CNN in low-dimensional datasets, while also proving to be essential in scaling Wormhole beyond 3D data.
 Fig.
@@
 While not trained on any ’Bag’ point cloud, their encodings still largely agree with true OT distance, producing a Pearson correlation of $\rho=0.995$ and label accuracy of $0.97$.
-For each dataset, we quantitatively benchmarked Wormhole by measuring how well the learned embeddings matched the true OT distance and ground-truth labels.
-Since the full Wasserstein distance matrix is unattainable, for each dataset we randomly sampled $10$ instances of $n=128$ point clouds and compared the $n\times n$ OT distance matrix to the embedding Euclidean distance.
-All the datasets are labeled, allowing us to measure the classification accuracy based on held-out test-sets for each.
-We benchmarked our Wormhole against DiffusionEMD and Deep Wasserstein Embedding (DWE), both sharing Wormhole’s goal of finding OT preserving embeddings.
-We also compare against a constructed baseline approach we dub Fréchet [48] in which we use the closed-form $W_{2}$ distance for Gaussian distributions.
-For each point cloud, we calculated the mean ($\mu$) and covariance matrix ($\Sigma$) of its samples and use the following formula to approximate the OT distance between each pair of point clouds:
-$W_{2}^{2}(X_{i},X_{j})\approx\|\mu_{i}-\mu_{j}\|^{2}+\text{Tr}\left(\Sigma_{i}+\Sigma_{j}-2\sqrt{\Sigma_{i}\Sigma_{j}}\right)$ (10)
-The following sections present vignettes of Wormhole from several dataset, and results are summarized in Table 1.
-Wormhole outperforms all competing approaches and uniquely scales to massive & high-dimensional datasets such as the the 256,058-sample, 254-dimensional MERFISH Cell Niches cohort.
-We specifically highlight the improvement of Wormhole over DWE, which similarly learns a neural network parametric embedding.
-The transformer backbone of Wormhole proves to be superior to DWE’s CNN mode in low-dimensional dataset, while also essential in scaling Wormhole beyond 3D data.
 6.1 MNIST (Touchstone)
@@
 We converted each sample to a point cloud by thresholding and selecting the set of pixel coordinates that passed the threshold (Figure 3a).
-While MNIST is considered small by modern machine-learning standards, straightforward computation of all $\binom{70000}{2}=2.45\cdot 10^{9}$ pairwise Wasserstein distances using current OT solvers is far from computationally feasible and would require a week of JAX-OTT [49] on an 80GB GPU (Figure 2).
+While MNIST is considered small by modern machine-learning standards, straightforward computation of all $\binom{70000}{2}=2.45\cdot 10^{9}$ pairwise Wasserstein distances using current OT solvers is far from computationally feasible and would require a week of JAX-OTT (Cuturi et al., 2022) on an 80GB GPU (Figure 2).
 By contrast, training Wormhole with default parameters (Supplementary Section 11) requires a small fraction of OT computations (approximately a $2000$-fold reduction) compared to the naive approach, with a total runtime of only 10 min.
@@
 Wormhole interpolation between two point clouds within the same class (top) and across different classes (bottom).
-Intermediate point clouds are linear interpolations between the embedding of each sample.
+Intermediate point clouds are decodings of linear interpolations between embeddings.
 As a more challenging test case, we assessed the ability of the Wormhole decoder to interpret the embeddings and solve variational Wasserstein problems.
@@
 Interpolating between distributions in Wasserstein space involves solving a weighted version of the barycenter problem as many times as there are intermediate samples.
-With Wormhole, we can leverage the Euclidean approximation of its latent space and directly decode linear interpolations between embeddings (Figure 5b).
-Despite the fact that interpolated embeddings are outside of the Wormhole training distribution, their decodings are valid and reliably visualize point cloud transitions.
-6.4 ShapeNet (Gromov-
-Wasserstein Wormhole)
-Gromov-Wasserstein (GW) [50] is an extension of OT to distributions in different spaces that seeks the most isometric map between point clouds. Given two point clouds, their GW distance is:
+With Wormhole, we can leverage the Euclidean approximation of its latent space and directly decode linear interpolations of embeddings (Figure 5b).
+Despite interpolated embeddings lying outside of Wormhole training distribution, their decodings reliably visualize point cloud transitions.
+6.4 ShapeNet (Gromov-Wasserstein Wormhole)
+Gromov-Wasserstein (GW) (Mémoli, 2011) is an extension of OT which seeks the most isometric map between point clouds. Given two point clouds, their GW distance is:
 $GW(X,Y)=\min_{P\in U(\nu,\mu)}\sum_{i,j,i^{\prime},j^{\prime}}\lVert d^{X}_{i,i^{\prime}}-d^{Y}_{j,j^{\prime}}\rVert^{2}P_{i,j}P_{i^{\prime},j^{\prime}}$ (11)
 Where $d^{X}_{i,i^{\prime}}$ is the distance between $X_{i}$ and $X_{i^{\prime}}$, and similarly for $d^{Y}_{j,j^{\prime}}$.
-This problem is quadratic and thus, it is significantly more computationally intensive than standard OT [2].
-GW has been a promising tool for aligning datasets from different modalities [51], mapping between graphs [52] and quantifying distances between 3D shapes with arbitrary orientation [53].
-However, the utility of GW is curtailed by its complexity, motivating our Wormhole approach.
+This problem is quadratic is significantly more computationally intensive than standard OT (Peyré et al., 2019).
+Unlike standard OT, GW can align datasets from different modalities (Demetci et al., 2020), map between graphs (Xu et al., 2019) and quantify distances between 3D shapes with arbitrary orientation (Mémoli, 2009).
 ShapeNet is a CAD-based dataset of 3D point clouds consisting of 16 different object classes.
 To construct the conditions where GW is usually applied, we randomly rotated each sample in 3D dimensions (Figure 6a).
+Standard OT is blind to structure when orientation is unknown, making rotation-invariant GW necessary.
 We then trained Wormhole to encode the rotated point clouds such that distances between embeddings retain GW distances.
-Although the training objective was radically different, Wormhole generalized to GW and learned accurate embeddings (Table 1).
-In addition, the encodings were reflective of object classes despite input coordinates being randomly rotated, implying that GW Wormhole was able to produce rotation invariant embeddings (Figure 6b).
 Fig.
@@
 Despite their random orientations, Wormhole was able to encode the ShapeNet point clouds and represent their class, as well as recover GW distances (Table 1).
+Although the training objective was radically different, Wormhole generalized to GW and learned accurate embeddings (Table 1).
+In addition, the encodings were reflective of object classes despite input coordinates being randomly rotated, implying that GW Wormhole was able to produce rotation invariant embeddings (Figure 6b).
 6.5 Cellular Niches (High-Dimensional Data)
 Spatially resolved transcriptomics (ST) are a set of technologies which capture both RNA information and the spatial context of cells to gain insight into tissue organization.
-OT has been frequently applied [54, 55] to analyze results of ST and unearth the relationship between a cell’s phenotype and its microenvironment.
-For example, Spatial Optimal Transport (SPOT) [56] models the niche of each cell as the point cloud from the gene expression of every cell in its environment, and defines differences between niches using Wasserstein.
+OT has been frequently applied to analyze results of ST and unearth the relationship between a cell’s phenotype and its microenvironment (Cang & Nie, 2020; Haviv et al., 2024).
+For example, Spatial Optimal Transport (SPOT) (Mani et al., 2022) models the niche of each cell as the point cloud from the gene expression of every cell in its environment, and defines differences between niches using Wasserstein.
 While OT is an alluring metric to quantify cellular niches, SPOT cannot scale to ST datasets of typical size.
 As experimental technologies mature, atlas-level spatial assays continue to grow in ubiquity and size.
-One instance is the $254$-gene MERFISH dataset of the motor cortex [44], which consists of 256,058 cells across 64 individual slices.
+For example, the $254$-gene MERFISH dataset of the motor cortex (Zhang et al., 2021), which consists of 256,058 cells across 64 slices.
 We summarize the $50\mu m$ radius around each cell in the MERFISH dataset to construct niche point clouds in $254$ dimensions.
@@
 While low-dimensional datasets can be manipulated to fit within their frameworks, the curse of dimensionality prevents adequate voxelization of this Cell Niches dataset.
-As commercial technologies capture more genes [57], the need for a method that can scale to higher dimensions, such as Wormhole, becomes more acute.
-Parvalbumin-positive (Pvalb) neurons are one of four major groups of inhibitory neuron, and are crucial for the regulation of neural circuitry.
-They are sparse in the motor cortex, comprising only $3\%$ of the dataset.
-To understand their spatial organization, it is therefore critical to study the entire cohort and integrate insights from multiple slices.
+As commercial technologies capture more genes (Marco Salas et al., 2023), the need for a method that can scale to higher dimensions, such as Wormhole, becomes more acute.
 Fig.
@@
 Wormhole embeddings reveal the cortical arrangement of these cells, as the first DC corresponds to neuronal depth.
+Parvalbumin-positive (Pvalb) interneurons are one the major groups of inhibitory neuron in the cortex, and are crucial for the regulation of neural circuitry (Sohal et al., 2009).
+They are sparse in the motor cortex, comprising only $3\%$ of the dataset.
+To understand their spatial organization, it is therefore critical to study the entire cohort and integrate insights from multiple tissue slices.
 After training Wormhole on the complete datasets (Table 1), we analyze the embeddings of the niches of Pvalb neurons.
-The first diffusion component (DC) [58] of Wormhole encodings aligns with their cortical depth across all slices within the dataset (Figure 7a-b).
+The first diffusion component (DC) (Coifman & Lafon, 2006) of Wormhole encodings aligns with their cortical depth across all slices within the dataset (Figure 7a-b).
 Despite the size and dimensionality of this dataset, Wormhole can still identify key biological signals across the cohort.
 7 Discussion
-We have introduced an approach, Wasserstein Wormhole, for embedding point clouds such that OT distances are preserved and can be approximated efficiently.
-Although OT is a popular approach, applying it to cohorts of distributions requires costly computation that scales quadratically with cohort size.
-Wormhole expands the functionality of OT to large datasets.
+We have introduced Wasserstein Wormhole, an approach for embedding point clouds such that OT distances are preserved and can be approximated efficiently.
+Although OT is ubiquitous, applying it to cohorts of distributions requires costly computations and scales quadratically with cohort size.
+Wormhole expands OT to large datasets.
 Encoding based on OT distances is theoretically challenged by the fact that Wasserstein is a non-Euclidean metric.
 To alleviate this concern, we derived theoretical lower and upper bounds that delimit how far non-Euclidean distance matrices can be from the embeddings.
-Our bounds grow smaller and tighter as the distance matrix approaches Euclidean.
 We further describe an algorithm based on projected gradient descent which is guaranteed to find an optimal embedding.
 On examples where the full Wasserstein distance matrix is tractable, Wormhole falls within the bounds and is close to the optimal error.
-We applied our approach to six different settings, ranging from 2D point clouds to massive spatial transcriptomics datasets with hundreds of genes.
+We applied our approach to eight different settings, ranging from 2D point clouds to massive spatial transcriptomics datasets with hundreds of genes.
 Across all settings, Wormhole embeddings recapitulated the Wasserstein space, accurately encoded key aspects of the point clouds and outperformed competing approaches.
@@
 Training Wasserstein Wormhole never exceeded 30 min on a single GPU, even on the 2048 points-per-sample ModelNet40 & ShapeNet datasets.
-Our algorithm is implemented in JAX [59] and integrates with OT tools (OTT-JAX) [49] for efficient Wasserstein distance calculations.
-Wormhole can be easily applied to any point cloud dataset via a pip installable package [^1] See tutorial in github.com/dpeerlab/WassersteinWormhole.
+Our algorithm is implemented in JAX (Bradbury et al., 2021) and integrates with OT tools (OTT-JAX) (Cuturi et al., 2022) for efficient Wasserstein distance calculations.
+Wormhole can be easily applied to any point cloud dataset via a pip installable package.
 Several methods have been proposed to learn approximations for the Wasserstein distance, including those based on neural networks.
@@
 We envision several future directions for this work.
-The complexity of transformers is quadratic in input size, limiting their scalability to point clouds with more than a few thousand samples.
-Alternatively, recently proposed models [60] can efficiently learn from point clouds by leveraging local geometry.
-Fitting those models within the Wormhole framework could produce a more versatile approach.
+Currently, Wormhole is only equipped to handle uniform point-clouds, where each point is equally weighted.
+Future work can extend the Attention mechanism to include arbitrarily weighted point-clouds, to reflect more general instances of empirical distributions and applications of OT.
 Wasserstein is not only used to define a distance between distributions, but also to find the optimal mapping across them.
@@
 This could possibly be achieved by exploiting the differentiability of Wormhole embeddings with respect to the point clouds.
+Acknowledgements
+This work was supported by National Cancer Institute grants P30 CA08748 and U54 CA209975. Dana Pe’er is an Investigator at the Howard Hughes Medical Institute.
+Impact Statement
+Wasserstein Wormhole is a novel method to scale computations of OT distances to large cohorts. We currently do not foresee any adverse effects of our work